# What’s the Weight?

### v1

@@ 3.4 Estimation
To estimate the ACD based on the three formulas presented above (OM, IPW1, and IPW2), we propose the following. For OM, define the functional
$\mu(a)=\mathbb{E}\left[\mathbb{E}[Y\mid A=a,S=1,X]\frac{\Pr(S=1)}{\Pr(S=1\mid X)}\mid S=1\right].$
We denote the estimator for the expectation of the outcome, $\mathbb{E}[Y\mid A=a,S=1,X]$, by $\hat{g}_{a}(X)$. In practice, $\hat{g}_{a}(X)$ can be estimated, for example, using a standard parametric outcome model with a finite-dimensional parameter vector. Our outcome model-based estimator for $\mu(a)$ is given by
$\hat{\mu}_{\rm OM}(a)=\frac{1}{n}\sum_{i=1}^{n}\hat{g}_{a}(X_{i})\frac{\Pr(S_{i}=1)}{\Pr(S_{i}=1\mid X_{i})}.$ (5)
[…]
$\hat{\mu}_{\rm IPW1}(a)$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{I(A_{i}=a)Y_{i}Pr(S_{i}=1)}{\hat{e}^{w}_{a}(X_{i})Pr(S_{i}=1\mid A_{i}=a,X_{i})}\quad{\rm or}$ (6)
$\hat{\mu}_{\rm IPW2}(a)$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{I(\mathcal{A}=a)Y_{i}Pr(S_{i}=1)}{\hat{e}_{a}(X_{i})Pr(S_{i}=1\mid X_{i})}.$ (7)
[…]
@@ 3.4.1 Estimation of Uncertain Sample Selection or Non-Probability Samples
[…]
@@ 3.5 Inference
$\phi(Y,X,\theta,\mathcal{P})=M^{-1}\psi[Y,X,\theta(\mathcal{P})],$
where
@@
is the Jacobian matrix of the estimating equations (Huber, 1996). Using numerical optimization, the covariance matrix of the estimated parameters, $\hat{V}(\hat{\theta})$, including the ACD estimator, is given by
$\hat{V}(\hat{\theta})=\sum_{i=1}^{n}\phi_{i}(Y,X,\hat{\theta},\mathcal{P})\phi_{i}(Y,X,\hat{\theta},\mathcal{P})^{\prime},$ (8)
where $\phi_{i}(Y,X,\hat{\theta},\mathcal{P})$ is the empirical evaluation of the $m\times 1$ vector of the $i$th observation’s values of the influence function for $\hat{\theta}$ by replacing the unknown parameters with sample estimates.
@@ 3.5.1 Inference in Stratified, Clustered Multi-Stage Survey Samples
NHANES employs a complex stratified, clustered sampling design. While bias incurred from the differential sample selection probabilities (encoded in the survey weights) is accounted for in the estimation procedure, inference is affected by the fact that individuals within strata and primary sampling units (PSUs) are correlated. To account for this, we modify this linearized variance estimator, known as the robust or sandwich estimator outside of survey-based analysis, by summing the contributions of the $i$th individual $(i=1,\ldots,n_{j})$, nested in the $j$th primary sampling unit (PSU; $j=1,\ldots,J_{k}$), nested in the $k$th sampling stratum $(k=1,\ldots,K)$. Thus, the estimated covariance matrix for our parameters is expressed in terms of the variability of the between PSU-level sums of the first order Taylor approximations within the sampling strata. This becomes
$\sum_{k=1}^{K}\frac{J_{k}}{J_{k}-1}\sum_{j=1}^{J_{k}}\left\{\phi_{\cdot jk}\left(y,\hat{\theta},\mathcal{P}\right)-\bar{\phi}_{\cdot\cdot k}\left(y,\hat{\theta},\mathcal{P}\right)\right\}\left\{\phi_{\cdot jk}\left(y,\hat{\theta},\mathcal{P}\right)-\bar{\phi}_{\cdot\cdot k}\left(y,\hat{\theta},\mathcal{P}\right)\right\}^{\prime}$ (9)
where $K$ and $J_{k}$ are the number of strata and number of PSUs within the $k$th stratum, respectively, $\phi_{\cdot jk}(y,\hat{\theta},\mathcal{P})=\sum_{i=1}^{n_{j}}\phi_{ijk}(y,\hat{\theta},\mathcal{P})$ is the sum of the influence function contributions among individuals in the $j$th cluster, and $\bar{\phi}_{\cdot\cdot k}(y,\hat{\theta},\mathcal{P})=\frac{1}{J_{k}}\sum_{j=1}^{J_{k}}\phi_{\cdot jk}$ is the “average” influence function value within the $k$th stratum. Software for these methods can be found in the R package svycdiff at https://github.com/salernos/svycdiff, in addition to the code and data necessary to reproduce the simulation studies and data analysis described below.
@@ 4 Simulation Studies
We conduct a series of simulation studies to better understand the finite sample performance of our proposed estimators, as compared to approaches typically used in practice. Based on the simulation studies conducted in DuGoff et al. (2014) and Ridgeway et al. (2015), we extend this work in several important ways. First, we have our proposed estimators, which represent novel alternatives to various combinations of weighting schemes outlined in Table 1. Second, we have varied the set of associations in our data to systematically include or exclude associations between $A$, $X$, and $S$ in our data generating mechanism to understand how certain relationships impact our estimation and inference. The purpose of the overall simulation study is to quantify the bias and variance associated with the use of our proposed approach, as compared to the various weighting schemes which may be currently employed.
@@ 4.1 Data Generation
As in DuGoff et al. (2014) and Ridgeway et al. (2015), let $X$ denote a normally-distributed covariate which relates to our outcome, $Y$, race, $A$, and sample selection indicator, $S$. We generate population-level data with $X\sim\mathcal{N}(1,1)$ for a super population of $N=100,000$ individuals from which we repeatedly take samples. We then generate the remaining data from three models:
  1. 1.
    Propensity Model: $\Pr(A\mid X)={\rm logit}^{-1}(\tau_{0}+\tau_{X}X)$,
  2. 2.
    Sample Selection Model: $\Pr(S\mid A,X)={\rm logit}^{-1}(\beta_{0}+\beta_{A}A+\beta_{X}X+\varepsilon_{S});\ \varepsilon_{S}\sim\mathcal{N}(0,0.1)$,
  3. 3.
    Outcome Model: $f(Y\mid A,X)$,
where ${\rm logit}^{-1}(u)=[1+\exp\{-u\}]^{-1}$ is the inverse logit function and we add a small amount of noise to the generation of the sample selection indicator to model this variable in our later procedure. From these models, we simulate the binary variable of interest, $A\mid X\sim\text{Bin}[N,\Pr(A\mid X)]$, the sample selection indicator, $S\mid A,X\sim\text{Bin}[N,\Pr(S\mid A,X)]$, and sample selection weights, $\omega_{S}=\Pr^{-1}(S\mid A,X)$. In simulating the outcome $Y$, we consider a heterogeneous treatment effect which varies with $X$, such that
$Y=\gamma_{0}+\gamma_{X}X+\gamma_{A}A+\gamma_{AX}AX+\varepsilon_{O}\$
where $\varepsilon_{O}\sim\mathcal{N}(0,1)$. We consider eight simulation settings, over which we vary the coefficients $\tau_{X},\beta_{A}$, and $\beta_{X}$, corresponding to the parameters in the true propensity and selection models (see Figure 2a-b), where each setting corresponds to a progressively more complex relationship among the variables. Across all simulation settings, we fixed $\tau_{0}=-1$, $\beta_{0}=-4.5$, $\gamma_{0}=1$, $\gamma_{A}=1$, $\gamma_{X}=1$, and $\gamma_{AX}=0.1$. Within each setting, we drew $200$ sub-samples based on our simulated sample selection indicator, $S$. We fit the various combinations of approaches listed in Figure 2. Our target quantity of interest was the ACD. We compared the percent bias, mean squared error (MSE), and coverage probability at a significance level of 0.05 for each method.

### v2

@@ 3.4 Estimation
To estimate the ACD based on the three formulas presented above (OM, IPW1, and IPW2), we propose the following. Let $\hat{g}_{a}(\boldsymbol{X})$ be an estimator of $g_{a}(\boldsymbol{X})=\mathbb{E}_{\mathcal{Q}}(Y\mid A=a,\boldsymbol{X})$. In practice, $\hat{g}_{a}(\boldsymbol{X})$ can be estimated, for example, using a parametric outcome model with a finite-dimensional parameter vector. Our outcome model-based estimator for $\mu(a)$ represented in Equation (1) is
$\hat{\mu}_{\rm OM}(a)=\frac{1}{n}\sum_{i=1}^{n}w(\boldsymbol{X}_{i})\hat{g}_{a}(\boldsymbol{X}_{i}).$ (5)
When the outcome model is correctly specified, $\hat{g}_{a}(\boldsymbol{X})\overset{p}{\rightarrow}g_{a}(\boldsymbol{X})$, leading to the result that $\hat{\mu}_{\rm OM}(a)\overset{p}{\rightarrow}\mu(a)$. Alternatively, IPW1 and IPW2 rely on inverse probability (IP) weighting to control for confounding (Horvitz and Thompson, 1952). Let $\hat{e}^{\mathcal{Q}}_{a}(\boldsymbol{X})$ denote the within-sample propensity model, which provides an estimator for $e_{a}^{\mathcal{Q}}(\boldsymbol{X})$. Similarly, let $\hat{e}^{\mathcal{P}}_{a}(\boldsymbol{X})$ denote the survey-weighted estimator of $e_{a}^{\mathcal{P}}(\boldsymbol{X})$. These are both estimable, as we assume a probability sample with known survey weights. A common approach may be to model the propensity via logistic regression, though many non-parametric approaches have been proposed (Lee et al., 2010). Depending on how we factorize the joint probability of sample selection and group membership, we can write the IPW estimators as
$\hat{\mu}_{\rm IPW1}(a)$ $=\frac{1}{n}\sum_{i=1}^{n}w_{a}(\boldsymbol{X}_{i})\frac{\mathbb{I}(A_{i}=a)Y_{i}}{\hat{e}^{\mathcal{P}}_{a}(\boldsymbol{X}_{i})},\quad{\rm or}$ (6)
$\hat{\mu}_{\rm IPW2}(a)$ $=\frac{1}{n}\sum_{i=1}^{n}w(\boldsymbol{X}_{i})\frac{\mathbb{I}(A_{i}=a)Y_{i}}{\hat{e}^{\mathcal{Q}}_{a}(\boldsymbol{X}_{i})}.$ (7)
If the propensity model is correct, that is $\hat{e}^{\mathcal{P}}_{a}(\boldsymbol{X})\pi_{a}(\boldsymbol{X})\overset{p}{\rightarrow}\Pr(S=1,A=a\mid\boldsymbol{X})$ and $\hat{e}^{\mathcal{Q}}_{a}(\boldsymbol{X})\pi(\boldsymbol{X})\overset{p}{\rightarrow}\Pr(S=1,A=a\mid\boldsymbol{X})$, then $\hat{\mu}_{\rm IPW1}(a)$ and $\hat{\mu}_{\rm IPW2}(a)$ are consistent.
3.4.1 Augmented Inverse Probability Weighted Estimator
Following Dahabreh et al. (2020), we can also derive the efficient influence function (EIF) for our setting. Specifically, the EIF is the canonical gradient for the parameter $\mu(a)$. Here, the EIF for $\mu(a)$ is given by
$\phi_{a}(Y,A,\boldsymbol{X})=\frac{\Pr(S=1)}{\Pr(S=1\mid\boldsymbol{X})}\left\{\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q}}_{a}(\boldsymbol{X})}\left[Y-g_{a}(\boldsymbol{X})\right]+g_{a}(\boldsymbol{X})\right\}-\mu(a)=w(\boldsymbol{X})\left\{\frac{\mathbb{I}(A=a)}{e^{\mathcal{Q}}_{a}(\boldsymbol{X})}\left[Y-g_{a}(\boldsymbol{X})\right]+g_{a}(\boldsymbol{X})\right\}-\mu(a).$
This also provides us the form of the doubly robust estimator, given by
$\hat{\mu}_{DR}(a)=\frac{1}{n}\sum_{i=1}^{n}w(\boldsymbol{X}_{i})\left\{\hat{g}_{a}(\boldsymbol{X}_{i})+\frac{\mathbb{I}(A_{i}=a)}{\hat{e}^{\mathcal{Q}}_{a}(\boldsymbol{X}_{i})}\left[Y_{i}-\hat{g}_{a}(\boldsymbol{X}_{i})\right]\right\},$ (8)
where $\hat{g}_{a}(\boldsymbol{X}_{i})$ is an estimator for $g_{a}(\boldsymbol{X})$ and $\hat{e}^{\mathcal{Q}}_{a}(\boldsymbol{X}_{i})$ is an estimator for the within-sample propensity score. For the ACD comparing groups 1 and 0, the final estimator is given by
$\widehat{ACD}_{DR}=\hat{\mu}_{DR}(1)-\hat{\mu}_{DR}(0).$
We provide the full derivation and proof in Appendix A.3.
@@ 3.4.2 Estimation of Uncertain Sample Selection or Non-Probability Samples
Throughout this work, we are motivated by settings where the sample arises from a known probability survey design. That is, we know $\pi_{a}(\boldsymbol{X})=\Pr(S=1\mid A=a,\boldsymbol{X})$. In practice, however, the sample selection mechanism may be unknown. For example, in non-probability samples, such as those arising from the study of electronic health records, we must estimate $\pi_{a}(\boldsymbol{X})$ with an assumed model ((e.g., through beta regression, as recommended by Elliot, 2009). In these settings, we replace $\pi_{a}(\boldsymbol{X})$ by a consistent estimator, $\hat{\pi}_{a}(\boldsymbol{X})$. Further, we can replace $\pi(\boldsymbol{X})=\Pr(S=1\mid\boldsymbol{X})$, the sample selection probability marginalized over $A$, by $\hat{\pi}(\boldsymbol{X})$, and $\Pr(S=1)$ by its estimate, $\hat{\pi}=n/N$. Here, our asymptotic results hold if the propensity and selection models are both correctly specified, then we have that $\hat{e}^{\mathcal{P}}_{a}(\boldsymbol{X})\hat{\pi}_{a}(\boldsymbol{X})\overset{p}{\rightarrow}\Pr(S=1,A=a\mid\boldsymbol{X})$ and $\hat{e}^{\mathcal{Q}}_{a}(\boldsymbol{X})\hat{\pi}(\boldsymbol{X})\overset{p}{\rightarrow}\Pr(S=1,A=a\mid\boldsymbol{X})$, leading to $\hat{\mu}(a)$ being consistent estimators of $\mu(a)$. As consistency depends on correctly specifying the selection model, alternative approaches, such as simplex regression or nonparametric methods, may provide more robust estimates (Kundu et al., 2023; Salvatore et al., 2024). If selection depends on additional, unmeasured covariates, $U$, that do not affect the group membership or the outcome, then our procedure remains unbiased, but we may underestimate the analytic standard errors. For inference, we must account for this additional step to obtain correct standard errors (see below).
@@ 3.5 Inference
$\sum_{i=1}^{n}\psi(O_{i};\boldsymbol{\theta})=0,$
where $\psi(O;\boldsymbol{\theta})$ stacks the estimating functions for the nuisance models (e.g., outcome and/or propensity, and selection where applicable; see Section 3.4.2) and the target components, and $\boldsymbol{\theta}=\boldsymbol{\theta}(\mathcal{Q})$ denotes the probability limit under $\mathcal{Q}$. Under standard regularity conditions, $\hat{\boldsymbol{\theta}}$ is asymptotically linear,
$\hat{\boldsymbol{\theta}}-\boldsymbol{\theta}(\mathcal{Q})=\frac{1}{n}\sum_{i=1}^{n}\phi(O_{i})+o_{p}(n^{-1/2}),$
with influence function
$\phi^{\mathcal{Q}}(O)=-M^{-1}\psi[O;\boldsymbol{\theta}(\mathcal{Q})];\quad M=-\mathbb{E}_{\mathcal{Q}}[\nabla_{\theta}\psi(O;\boldsymbol{\theta})]_{\boldsymbol{\theta}=\boldsymbol{\theta}(\mathcal{Q})}.$
Here, $M$ is the Jacobian matrix of the estimating equations (Huber, 1996). The population asymptotic covariance matrix is then given by
$V(\theta)=\frac{1}{n}\Sigma(\theta);\quad\Sigma(\theta)=\mathbb{E}_{\mathcal{Q}}\left[\left(\phi^{\mathcal{Q}}\right)\left(\phi^{\mathcal{Q}}\right)^{\top}\right]=\left[M(\theta)^{-1}\right]B(\theta)\left[M(\theta)^{-1}\right]^{\top};\quad B(\theta)=\mathbb{E}_{\mathcal{Q}}\left[\psi\psi^{\top}\right].$
@@
$\hat{\Sigma}_{n}=\frac{1}{n}\sum_{i=1}^{n}\hat{\phi}_{i}\hat{\phi}_{i}^{\top}=\left[\hat{M}^{-1}\right]\left\{\frac{1}{n}\sum_{i=1}^{n}\hat{\psi}_{i}\hat{\psi}_{i}^{\top}\right\}\left[\hat{M}^{-1}\right]^{\top}.$
Consequently, the estimated covariance matrix of $\hat{\theta}$ is
$\hat{V}(\hat{\theta})=\frac{1}{n}\hat{\Sigma}_{n}=\frac{1}{n^{2}}\sum_{i=1}^{n}\hat{\phi}_{i}\hat{\phi}_{i}^{\top}=\frac{1}{n}\left[\hat{M}^{-1}\right]\left\{\frac{1}{n}\sum_{i=1}^{n}\hat{\psi}_{i}\hat{\psi}_{i}^{\top}\right\}\left[\hat{M}^{-1}\right]^{\top}.$
[…]
@@ 3.5.1 Inference in Stratified, Clustered Multi-Stage Survey Samples
[…]
$\hat{V}(\hat{\theta})=\frac{1}{n^{2}}\sum_{k=1}^{K}\frac{J_{k}}{J_{k}-1}\sum_{j=1}^{J_{k}}\left(\hat{U}_{jk}-\bar{U}_{k}\right)\left(\hat{U}_{jk}-\bar{U}_{k}\right)^{\top}.$ (9)
3.6 Software
Software for these methods can be found in the R package svycdiff, available through the Comprehensive R Archive Network (CRAN) at cran.r-project.org/web/packages/svycdiff, or in a development version on GitHub at github.com/salernos/svycdiff. As our proposed approach uses an estimating equations framework, the svycdiff accommodates various outcome types, including continuous, binary, and count outcomes. This includes the main functions for estimation and inference, as well as a function for simulating population-level data under a wide range of scenarios. Moreover, these accommodate the setting described in Section 3.4.2. Lastly, the codes necessary to reproduce our later simulations and data analysis are also available at the GitHub link.
@@ 4 Simulation Studies
We conduct a series of simulations based on the setup of Austin et al. (2018), and subsequently Zeng et al. (2025), to understand the finite sample performance of our proposed estimators compared to those typically used in practice. We systematically vary the associations between $A$, $\boldsymbol{X}$, and $S$ in our data-generating mechanism to understand how these relationships impact estimation and inference. We quantify the bias and variance of our proposed approach compared to existing approaches.
@@ 4.1 Data Generation
We first generate population-level data for $N=1,000,000$ individuals. We organize the individuals into 10 strata $(j=1,\ldots,10)$ that are further subdivided into 20 clusters, so that there are 200 unique clusters $(k=1,\ldots,200)$ and 5,000 individuals per cluster $(i=1,\ldots,5,000)$. For each individual, we generate two baseline covariates $(l=1,2)$ from normal distributions, allowing the means of these covariates to vary at both the stratum and cluster levels. That is, we have
$X_{l,ijk}\sim\mathcal{N}\left(\nu_{l,j}^{s}+\nu_{l,j}^{c},1\right);\quad\nu_{l,j}^{s}\sim\mathcal{N}\left(0,\sigma_{l}^{s}\right),\quad\nu_{l,j}^{c}\sim\mathcal{N}\left(0,\sigma_{l}^{c}\right),$
where $\nu_{l,j}^{s}$ and $\nu_{l,j}^{c}$ are stratum- and cluster-specific random effects for the $l$th covariate. We set $\sigma_{l}^{s}=0.35$ and $\sigma_{l}^{c}=0.15$ across all simulation settings. Note that $\boldsymbol{X}_{i}=\{X_{1,i},X_{2,i}\}$ relate to our outcome, $Y_{i}$, group variable of interest, $A_{i}$, and sample selection indicator, $S_{i}$. We then generate $A_{i}\in\{0,1\}$ from a logistic model, so the propensity model for $A_{i}=1$ is given by
$\Pr(A_{i}=1\mid\boldsymbol{X}_{i})={\rm logit}^{-1}\left[\tau_{0}+\tau_{A}\left(\tau_{X_{1}}X_{1,i}+\tau_{X_{2}}X_{2,i}+\tau_{X_{1}X_{2}}X_{1,i}X_{2,i}\right)\right],$
where ${\rm logit}^{-1}(u)=[1+\exp\{-u\}]^{-1}$ is the inverse logit function. Here, $\tau_{0}$ controls the overall group prevalence, $\tau_{A}$ reflects the strength of confounding and, correspondingly, the degree of overlap in the propensity score, $\tau_{X_{1}}$ and $\tau_{X_{2}}$ are coefficients for $X_{1}$ and $X_{2}$, respectively, and $\tau_{X_{1}X_{2}}$ an interaction effect between $X_{1}$ and $X_{2}$. For this main simulation, we fix $\tau_{0}=0$, $\tau_{A}=0.5$, and $\tau_{X_{1}X_{2}}=0$, and we vary $\tau_{X_{1}}$ and $\tau_{X_{2}}\in\{0,1\}$. We then simulate the $A_{i}$ from a Bernoulli distribution with $A_{i}\sim{\rm Bernoulli}[\Pr(A_{i}=1\mid\boldsymbol{X}_{i})]$. We next generate a survey weight for each individual, $\omega_{i}$, by specifying a logistic selection model, with
$\Pr(S_{i}=1\mid A_{i},\boldsymbol{X}_{i})={\rm logit}^{-1}(\beta_{0}+\beta_{A}A_{i}+\beta_{X_{1}}X_{1,i}+\beta_{X_{2}}X_{2,i}+\varepsilon_{S});\quad\varepsilon_{S}\sim\mathcal{N}(0,0.1),$
where $\beta_{0}$ controls the marginal sampling fraction, $\beta_{A}$, $\beta_{X_{1}}$, and $\beta_{X_{2}}$ are coefficients for $A$, $X_{1}$ and $X_{2}$, respectively, and $\varepsilon_{S}$ is small amount of random noise added to the generation of the survey weights, to support later sensitivity analyses. We fix $\beta_{0}=-8$ and vary $\beta_{A}$, $\beta_{X_{1}}$, and $\beta_{X_{2}}\in\{0,1\}$, and we set $\omega_{i}=\Pr(S_{i}=1\mid A_{i},\boldsymbol{X}_{i})^{-1}$. Lastly, we simulate continuous outcomes, $Y_{i}$, for each individual in the population as a linear combination of $A_{i}$, $\boldsymbol{X}_{i}$, and interaction terms. Specifically, we let
$Y_{i}=\alpha_{0}+\alpha_{X_{1}}X_{1,i}+\alpha_{X_{2}}X_{2,i}+A_{i}\left[\alpha_{A}+\alpha_{AX}\left(\alpha_{X_{1}}X_{1,i}+\alpha_{X_{2}}X_{2,i}\right)\right]+\varepsilon_{O};\quad\varepsilon_{O}\sim\mathcal{N}(0,1),$
where we set $\alpha_{0}=1$ and $\alpha_{X_{1}}=\alpha_{X_{2}}=1$ for the covariate coefficients. We further set $\alpha_{A}=1$ to reflect the true effect of $A$ on $Y$ and $\alpha_{AX}=0.1$ to induce a heterogeneous group/exposure effect which varies with $\boldsymbol{X}$.
For this main simulation study, we consider a fully factorial design where we vary $\tau_{X_{1}}=\tau_{X_{2}}\in\{0,1\}$, $\beta_{A}\in\{0,1\}$, and $\beta_{X_{1}}=\beta_{X_{2}}\in\{0,1\}$, yielding eight parameter settings, where each setting corresponds to a progressively more complex relationship among the variables: two with no confounding or selection bias, two with confounding bias but no selection bias, two with selection bias but no confounding bias, and two with both selection and confounding bias (see Table LABEL:tab:simulation_results).
Within each simulation setting, we draw 500 random samples from our generated population by repeatedly simulating sample selection indicators, $S_{i}$. As the focus of the current work is in setting of known survey weights, which can be stratified or multistage, we extract the probability of selection for the $i$th individual and generate $S_{i}$ from a Bernoulli distribution with $S_{i}\sim{\rm Bernoulli}[\Pr(S_{i}=1\mid A_{i},\boldsymbol{X}_{i})]$, retaining those observations with $S_{i}=1$ in each replicate. Note that this deviates slightly from a multistage stratified, clustered design. However, we opted for this simplification to focus on the more important aspect of the problem, namely that the final weights need to appropriately factorize the joint distribution of being sampled and belonging to a particular group. This holds over a range of sampling designs (see our sensitivity analyses below).
We compare our proposed (5) – (8) to the standard analytic approaches in Table 1. Across all methods where survey weights are considered, we apply the corresponding appropriate standard error calculations. For example, for our proposed approaches, we use the variance formula we have derived for stratified, clustered designs (9). For comparison methods such as survey-weighted regression, we use design-based standard errors (Lumley, 2011). The oracle estimator, which knows the full data generating process, serves as our benchmark throughout. Our target quantity of interest is the ACD. We compare the relative bias, average of the estimated standard error (ASE), Monte Carlo standard error (MCSE), mean squared error (MSE), and coverage probability at a significance level of 0.05 for each method.